OpenStudy (anonymous):

Find the cube roots of 8(cos 216 + i sin 216)

OpenStudy (anonymous):

Well $$ \Large{ 8(cos(216) + i \cdot sin(216)) = 8 \cdot e^{i \cdot 216} \\ \; \\ \sqrt[3]{8 \cdot e^{i \cdot 216}} = \sqrt[3]{8} \cdot \Big(e^{i \cdot 216} \Big)^\frac{1}{3} = 2 \cdot e^{\big(i\frac{216}{3}\big)} = \\ = 2 \cdot e^{i \cdot 72} = 2(cos(72) + i \cdot sin(72)) } $$

OpenStudy (anonymous):

and 2(cos(72)+i⋅sin(72)) is the cube root?

OpenStudy (anonymous):

\[2(\cos(72)+i \times(72))\]*

OpenStudy (anonymous):

where did the sin() disappear?

OpenStudy (anonymous):

oh sorry I forgot to write that

OpenStudy (anonymous):

but ye, that's the cube root =]

OpenStudy (anonymous):

awesome thanks so much

OpenStudy (anonymous):

You're welcome It is all clear, right?